OpenStudy (anonymous):

each leg of a 45 45 90 triangle has a length of 6 units what is the length of its hypotenuse? a 6 units b 12 units c 3 square root over 2 units d 6 square root over 2 units

OpenStudy (lgbasallote):

do pyhtagorean theorem. it says both the legs are 6 so \[c = \sqrt{a^2 + b^2}\] \[c = \sqrt{6^2 + 6^2}\] \[c = \sqrt{36 + 36}\] \[c = \sqrt{72}\] simplify that and yoou get your answer

OpenStudy (anonymous):

ok so its 6 units right

OpenStudy (lgbasallote):

nope

OpenStudy (anonymous):

\[c=\sqrt{72}\]\[c=\sqrt{36*2}\]

OpenStudy (lgbasallote):

^that

OpenStudy (anonymous):

you will pull out the 36 and it will turn into a 6 when you pull it out

OpenStudy (anonymous):

i got it wrong thanks for helping though